A body X, of mass 12Kg is placed 3metre above the ground, if another body Y, of mass 20kg moving with   velocity V, has a kinetic energy equal in value to the potential energy of X, calculate the magnitude of V.(g=10 m/s).


1
Expert's answer
2021-04-09T06:34:18-0400

The potential energy of body X is:


"E_p = m_xgh"

where "m_x = 12kg" is the mass of X, "h = 3m" is its height, and "g = 10m\/s^2" is the gravitational acceleration.

The kinetic energy of Y is:


"E_k = \dfrac{m_yv^2}{2}"

where "m = 20kg" is the mass of Y, and "v" is its speed. SInce


"E_p = E_k"

obtain:


"m_xgh = \dfrac{m_yv^2}{2}\\
v^2 = \dfrac{2ghm_x}{m_y}\\
v = \sqrt{\dfrac{2ghm_x}{m_y}}"

Substituting the values, obtain:


"v = \sqrt{\dfrac{2\cdot 10\cdot 3\cdot 12}{20}} = 6m/s"

Answer. 6 m/s.